OpenStudy (anonymous):

find the base b if the graph of y=b^x contains the given points (-3,.008)?

OpenStudy (anonymous):

0.008 = b^(-3) Ln(0.008) = -3Lnb Lnb = Ln(0.008)/-3

OpenStudy (anonymous):

Ln(b) = 1.609 b = e^1.609 = 4.99

OpenStudy (anonymous):

or b = 5

OpenStudy (anonymous):

Please click on Good Answer if I helped, I really want to get to the next level, thanks =)!